Youtube 6. Find an indefinite integral. x2(x + 2) da 7. Find an indefinite integral. cos²ix dx
Answer #1

Intzar take partial fraction: x²(x+2) = 20 tu unto 1 1 1 hence: - de dx= u2(x+2) . J 2x2 .. 4(x+2) 4x talle a man thente [ en21 at haluis Ans. I t kl {cxis tx] = zep (1x950 til the tuotsus senoje IS 2 *P(1x710+ (129)10) to re-271) Fi Rep (X®•2)+ f =feel free to ask if you have any doubt and if you liked my answer please give a thumbs up